OpenStudy (anonymous):

Find the polynomial equation represented with points at: (0,0) (3,0) and (4,3.2) I found the equation of f(x)=.2(x+2)^2 (x-2)^3 This is wrong...can you please explain where my mistake is?

OpenStudy (anonymous):

this is my work, f(x)=a(x+0)^2(x-3)^3 3.2=a(4+0)^2(4-3)^3 3.2=a(16)(1) a=.2 f(x)=.2(x+2)^2(x-2)^3 where am I going wrong?

OpenStudy (anonymous):

you have 3 points, should be a quadratic polynomial

OpenStudy (anonymous):

y intercepts at 0 and 3 tells you it looks like \(p(x)=ax(x-3)\) and solve for \(a\) by taking \(p(4)=a\times 4\times (4-3)=3.2\)

OpenStudy (anonymous):

I still don't get it...

OpenStudy (anonymous):

you have three points a quadratic looks like \(p(x)=ax^2+bx+c\) with three unknowns, namely \(a,b,c\) so if you have three points, you are specifying a quadratic polynomial (degree two ) not some higher degree one i am not sure where you got the exponent of 3 and 2 from in your answer

OpenStudy (anonymous):

the 3 and 2 are the double zero and triple zero

OpenStudy (anonymous):

you have the zeros are \(0\) and \(3\) which means it factors as \(x(x-3)\) except the leading coefficient does not have to be 1, so it is \[p(x)=ax(x-3)\]

OpenStudy (anonymous):

does it say that 0 is a zero with multiplicity two?

OpenStudy (anonymous):

it didn't say that above, but if so, no matter, you would be correct and it would be \[f(x)=ax^2(x-3)^3\]

OpenStudy (anonymous):

no, it showed me the graph and I had to find the equation... I would show you the graph, but I d'ont think that I can

OpenStudy (anonymous):

ok so lets imagine that it has a zero at 0 with multiplicity 2, (so it touches right?) and somehow you know that the zero at 3 has multiplicity 3 then your work is right, but i don't know how you changed the correct work to get a different answer

OpenStudy (anonymous):

\[f(4)=3.2=16a\] so \(a=0.2\) and your answer is \[f(x)=.2x^2(x-3)^3\]

OpenStudy (anonymous):

? lost

OpenStudy (anonymous):

you have the answer. you can multiply it out of you like, but if it factors as \(f(x)=0.2x^2(x-3)^3\) it will not factor some other way

OpenStudy (anonymous):

you already have it in factored form. you can leave it like that, or you can multiply out but you cannot multiply out, then factor again and get something other than what you started with

OpenStudy (anonymous):

ok, so pretend that there is a graph with points (-2,0) (1,0) and (-1,-4) it would be f(x)=a(x+2)^2(x-1)^3 right?

OpenStudy (anonymous):

yes, assuming that for some reason you knew the multiplicity of the zeros not sure how you know it though

OpenStudy (anonymous):

if I was looking at another graph... then if than equation was right, then f(x)=2(x+2)^2(x-1)^3 and then I would get the same points...

OpenStudy (anonymous):

but I'm not...